OpenStudy (anonymous):

free medal and fan. x^-2 y^2/x^4 y^-3 x=4 and y=-2

OpenStudy (anonymous):

Hi, can you help me with this?

OpenStudy (anonymous):

What should we do? just solve?

OpenStudy (anonymous):

Yes, evaluate the expression.

OpenStudy (anonymous):

you just want the solution?

OpenStudy (anonymous):

yes, explanation if possible.

OpenStudy (anonymous):

Ugh, I also have another problem, The volume of a sphere is found using the formula V=4/3 3.14r^3. The area of a circle is found using the formula a=3.14r^2. What is the ratio of the volume to the area?

OpenStudy (anonymous):

I got the ratio of the volume and area question. Just need help on the original.

OpenStudy (anonymous):

\[\large \frac{x^{-2} y^2}{x^4 y^{-3}}\]

OpenStudy (anonymous):

first thing to do is rewrite it as \[\large \frac{y^5}{x^6}\]

OpenStudy (anonymous):

correct, x=4 and y=-2

OpenStudy (anonymous):

then go ahead and put in \(x=4,y=-2\) and get \[\frac{(-2)^5}{4^6}\]

OpenStudy (anonymous):

i believe the "final answer" will be \[\large -\frac{1}{128}\]

OpenStudy (anonymous):

Ok, thank you so much! I was trying to do the math. I got the same answer.

OpenStudy (anonymous):

wait,what?